Please help with how you calculate directional derivative

In this question, given the function f(x,y,z,w)= [x^2y+zw, y^2z+xw, xyzyw] find the differential df and the directional derivative at [-1,0,1,2]




1
Expert's answer
2021-02-24T12:05:38-0500

"f(x,y,z,w) = (x^{2}y + zw,y^{2}z + xw, xyz, yw)"

"df = 2yxdx+2zydy + xydz + ydw"

"\cfrac{\partial f}{\partial l} = \cfrac{\partial f}{\partial x}cos\alpha + \cfrac{\partial f}{\partial y} cos\beta + \cfrac{\partial f}{\partial z}cos\gamma + \cfrac{\partial f}{\partial w} cos\theta"

Find the direction cosines:

"l = (-1;0;1;2)\\
l_0 = \cfrac{l}{|l|} = \cfrac{-i+k+2t}{\sqrt{1+0+1+4}} = -\cfrac{1}{\sqrt{6}}i+\cfrac{1}{\sqrt{6}}k + \cfrac{2}{\sqrt{6}}t"

so we have:

"cos \alpha = -\cfrac{1}{\sqrt{6}},cos\beta = 0, cos\gamma = \cfrac{1}{\sqrt{6}}, cos\theta = \cfrac{2}{\sqrt{6}}"

so directional derivative at "(-1;0;1;2)" :

"\cfrac{\partial f}{\partial l} = -\cfrac{2yx}{\sqrt{6}} + \cfrac{xy}{\sqrt{6}} + \cfrac{2y}{\sqrt{6}} = -\cfrac{yx+2y}{\sqrt{6}}"